OpenStudy (anonymous):

If cos x cos (pi/7) + sin x sin (pi/7) =-√2/2, then x can equal: more than one answer: a. 3pi/4+pi/7+2 n pi b. 7 pi/4 + pi/7 +2 n pi c. pi/4 + pi/7 +2n pi d. 5pi/4 + pi/7 +2 n pi

OpenStudy (binary3i):

\[\cos a \cos b- \sin a \sin b=\cos (a+b)\] use this. now solve.

OpenStudy (anonymous):

can you help more please

OpenStudy (binary3i):

\[\cos x \cos \pi/7 + sinx \sin \pi/7= \cos(x-\pi/7)= -\sqrt{2}/2=-1/\sqrt{2}\] \[\cos \theta=-1/\sqrt{2}.. then \theta = 3\pi/4 + 2n \pi =(x-\pi/7)\] now try to solve it.

OpenStudy (binary3i):

answer is a.